A transverse wave is incident on a boundary separating two media of acoustic impdances Z1 and Z2.
(a) Write down the equations of incident, reflected and transmitted waves. .
(b) Write down the boundary conditions.
(c) Obtain expressions for reflection and
transmission coefficients
1
Expert's answer
2019-05-17T11:19:19-0400

1) For example, consider the following incident wave with incident angle "\alpha", refracted with angle "\beta" and transmitted with angle "\gamma":


"p_i=p_0e^{ik_1\vec{r}}=p_0\text{exp}[ik_1(x\text{cos}(90^\circ-\alpha)+y\text{cos}\alpha],"

reflected wave:


"p_r=p_0R\text{exp}[ik_1(x\text{sin}\beta+y\text{sin}\beta]"

and transmitted wave


"p_t=p_0D\text{exp}[ik_2(x\text{sin}\gamma+y\text{cos}\gamma],"

where "k_1,\space k_2" - wavenumbers for the mediums.

2) The boundary conditions (y=0):


"p_0\text{exp}(ik_1x\text{sin}\alpha)+Rp_0\text{exp}(ik_1x\text{sin}\beta)=Dp_0\text{exp}(ik_2x\text{sin}\gamma)."

According to Snellius law:


"k_1\text{sin}\alpha=k_1\text{sin}\beta=k_2\text{sin}\gamma,"

then


"D=R+1."

The boundary conditions at x=0:


"\frac{\text{cos}\alpha}{c_1\rho_1}-R\frac{\text{cos}\beta}{c_1\rho_1}=D\frac{\text{cos}\gamma}{c_2\rho_2}."

3) The reflection and transmission coefficients are


"R=\frac{Z_1-Z_2}{Z_1+Z_2},"

"D=\frac{2Z_1}{Z_1+Z_2}."